# Is my Data in your AI Model? Membership Inference Test with Application to Face Images

### v3

@@ Abstract
keywords:
Audit, Fairness, Reliability, Membership Inference, MIA, MINT, Face Recognition
journal: Information Fusion
@@
[1]organization=Biometrics and Data Pattern Analytics Lab, Universidad Autonoma de Madrid, country=Spain
1 Introduction
@@
Nowadays, AI systems have become increasingly advanced and pervasive across different domains, collecting, analyzing, and processing vast amounts of data. These data often contain sensitive information [3] about individuals as well as copyright content [4], raising concerns about privacy [5, 6, 7] and unauthorized access [8, 9, 10]. Consequently, it results critical the fast development of novel tools to clarify what data has been used to train AI models, preventing therefore the use of non-authorized data and increasing the transparency and explainability of the AI models [11, 12, 13]. These tools aim to prevent learning frameworks from hiding the training data within the model’s parameters, thereby unveiling the opacity that they have traditionally relied upon. These considerations lead us to the main objective of the present study: checking if given data was used to train AI models [14], as summarized in Fig. 1.
@@
The article is structured as follows. Sec. 2 provides a revision of the state of the art in the scope of the article. Sec. 3 describes the main concepts of the proposed MINT, and the specific details for the task considered, Face Recognition. The databases, Face Recognition systems, and experimental protocol are described in Sec. 4. Results are discussed in Sec. 5, whereas the significance of Face Recognition in MINT and the challenges of developing this technology in real-world applications are discussed in Sec. 6 and 7. Discussion and conclusions are finally drawn in Sec. 8 and 9.
2 Related Works
@@
2.1 Membership Inference Attacks
@@
2.2 Reconstructing Training Data from AI Models
@@
2.3 Data Tracing
Methods described in previous sections can be seen as “passive methods” that rely on an already trained model without interfering with it. In contrast, the Data Tracing research direction falls under the category of “active methods”, where the training data is directly altered to influence the model [51]. Data tracing involves changing the training data to identify if it was used during training. Several studies have tried to tackled this challenge, demonstrating that with smart changes in the training data, it is feasible to detect its presence during inference [52, 53]. For instance, Radioactive Data [54] achieved this by inserting “radioactive marks” into the training data to detect if a database was used to train a model.
@@ 2.3 Data Tracing
3 Membership Inference Test (MINT): Proposed Approach
@@
3.1 Problem Statement and Terminology
@@
We hypothesize that an authorized auditor possesses access to model $M$, enabling him the acquisition of information regarding how $M$ processes data $d$. Alternatively, the auditor may possess information detailing how $M$ has processed data $d$, even in the absence of direct model access. This information comprises the generated Model Outcome $y=M(d|\textbf{w})$ and if possible also some intermediate results (e.g., activation maps of specific layers in a neural network). These intermediate outcomes $N(d|\textbf{w}^{\prime})$ provide insights about a subset of parameters $^{\prime}\subset\textbf{w}$ w. We define these intermediate outcomes as Auxiliary Auditable Data (AAD). The auditor does not need the full description of the model $M$ or the values of the complete set of parameters w to obtain the AAD.
@@
  • 5.
    Auxiliary Auditable Data $\textrm{AAD}=N(d|\textbf{w}^{\prime})$: intermediate outcomes of $M$ that result from processing an input data $d$ using a subset $\textbf{w}^{\prime}$ of the parameters w. The model outcome can be seen as the case where $^{\prime}=\textbf{w}$ w in which case $N(d|\textbf{w}^{\prime})=M(d|\textbf{w})$.
  • 6.
@@
3.2 MINT: Application to Face Recognition Models
In the previous Sec. 3.1, we have introduced the general concepts of our MINT proposal, which can be applied to any type of Audited Model $M$. Next, we detail the various components of the proposal for the application of MINT to the challenging task of Face Recognition.
@@ 3.2 MINT: Application to Face Recognition Models
  • 2.
    $\mathcal{E}$ consists of images from external Face Recognition databases. We could add images unrelated to faces in $\mathcal{E}$. However, we included images with similar characteristics to $\mathcal{D}$ in order to prevent the network from learning to detect database-dependent features of face images instead of focusing in the main task: database-independent detection of used or not for training. It is crucial that these images do not overlap with those in $\mathcal{D}$, i.e., $\mathcal{E}\notin\mathcal{D}$.
@@
3.3 Proposed MINT Models
@@ 4 Experimental Framework
4.1 Face Recognition Data and Models
@@
As external data ($\mathcal{E}$) to train and test the MINT model $T$ we use the following datasets: The IJB-C [69] (3.5K identities and 138K face images), FDDB [70] (5.2K face images), GANDiffFace [71, 72] (10K identities and 500K face images), and Adience [73] (2.2K identities 26.5K face images). As mentioned in Sec. 3.2, it is crucial to avoid overlap between $\mathcal{E}$ and $\mathcal{D}$, as this would lead to incorrect labels, compromising the training of MINT Models $T$. As noted by Schlett et al. [74], there is a significant number of duplicates across commonly used datasets. Although the data in $\mathcal{E}$ is sourced from evaluation datasets that strive to prevent such overlap, the necessary tests have been conducted to ensure that this issue does not occur.
4.2 Experimental Protocol
@@
  • 2.
    $4$ different databases as External Data $\mathcal{E}$. Three of them (IJB-C, FDDB, GANDiffFace) are used to extract samples for training the MINT Model, while the fourth one (Adience) is used for evaluation. We will refer to this choice of databases as the “Baseline Case”. Having either IJB-C or GANDiffFace, or both, in the training set is necessary to achieve balanced training (enough $\mathcal{E}$ samples). If we include only GANDiffFace from these two databases in the training set, the network might learn to recognize specific patterns from synthetic images instead of focusing on the MINT task. Therefore, we always include IJB-C in the training set. However, the remaining three databases are interchangeable, allowing evaluation with GANDiffFace or FDDB and utilization of Adience in training. We refer to these two cases as “Case A” and “Case B”, respectively. We present some results interchanging this evaluation database in our experiments. The Baseline case provides the most robust evaluation, as Case A involves evaluation with synthetic images, which may raise doubts due to the nature of these images, and Case B involves evaluation with only 5.2K images. Therefore, the majority of results will be presented under the “Baseline Case”.
  • 3.
@@ 4.2.1 Data Pre-processing
4.2.2 Auxiliary Auditable Data and Model Outcome
@@
4.2.3 MINT Models Training
With the previously obtained AAD and Model Outcome, we propose two MINT models. Depending on how we handle the acquired information, we propose two architectures (Sec. 3.3): a Vanilla MLP model and a CNN model. The CNN utilizes the activations of the convolutional blocks of $M$ as-is, while the Vanilla model, as previously explained, applies max pooling per channel to convert the activations of the convolutional blocks into a vector that can be processed by the MLP.
@@ 4.2.3 MINT Models Training
Both architectures are chosen after extensive experimentation. It was found that the architectures did not need to be overly complex, as performance did not improve significantly with more complex architectures. In the experimental section (Section 5.1), an experiment has been included to demonstrate the model’s performance versus the network complexity.
5 Experiments and Results
This section presents the results achieved by our proposed MINT approach. As it involves binary classification between $\mathcal{D}$ and $\mathcal{E}$, and considers a balanced evaluation, the baseline accuracy (i.e., random guess) stands at 50%. Additionally, we remind that all results are derived from samples in $\mathcal{E}$ that do not belong to the $\mathcal{E}$ databases used in training the MINT Model, as elaborated in Sec. 4.2.
@@ 5 Experiments and Results
  • 1.
    FR Model ($M$ in Fig. 3): These are the Face Recognition models presented in Sec. 4.1. The number corresponds to the one specified in that section.
  • 2.
    Model Outcome: Refers to the Output Embedding of the Face Recognition model in Sec. 3.2.
  • 3.
    Conv Layer X: Represents the AAD obtained from each of the four primary layers, as explained in Sec. 4.2.2
  • 4.
@@
  • 5.
    Training Data: Represents three possible scenarios based on the amount of data available for training our MINT Model $T$. The high scenario ($100$K) assumes a significant amount of data, with $50$K data $\mathcal{D}$ and $50$K external data $\mathcal{E}$. In the medium scenario ($50$K), there are $25$K data samples each for $\mathcal{D}$ and $\mathcal{E}$. The low scenario ($1$K) considers only $500$ data samples for both $\mathcal{D}$ and $\mathcal{E}$. To put these numbers in context, $50$K data samples from $\mathcal{D}$ represent only $1\%$ of the MS1Mv3 database used for training and $0.3\%$ of Glint360k.
5.1 Results
Table 1 shows the classification accuracy results for the $3$ Face Recognition models using different information to train the Vanilla MINT Model $T$ (AAD and Model Outcomes). All of this is done under a scenario of high data available for training. In these results, we can observe a superiority in accuracy when combining all the AAD information (Combination). In terms of the AAD it is important to highlight that in FR Model 1, Conv Layer #2 yields the best results after Conv Layer #1, while in FR Model 2 and 3, Conv Layer #4 follows with better accuracy. This might be attributed to the fact that ResNet architectures are based on “Skip Connections”, allowing information to traverse through the network from first layers to the last ones [76].
Table 1: Classification accuracy for various AAD and Model Outcomes using the Vanilla MINT Model. MINT model trained with $100K$ samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3).
@@ 5.1 Results
The results in this section pertain to the Baseline Case presented in Sec. 4.2, where the databases IJB-C, FDDB, and GANDiffFace are used in training the MINT Models, while Adience is used for evaluation. This is the case that provides a more robust evaluation (Sec. 4.2), and therefore the one mostly used throughout the experimental section. However, to demonstrate the robustness of the results, Table 2 showcases the same results as Table 1 with exchanged training and evaluation databases for the other two scenarios (see Sec. 4.2):
  • 1.
@@
As shown in Table 2, the outcomes are slightly superior for Case A and slightly inferior for Case B compared to the Baseline Case, yet the the conclusions remain unchanged (results are presented as Case A / Case B).
Table 2: Classification accuracy for various AAD and Model Outcomes using the Vanilla MINT Model. MINT model trained with $100K$ samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3). We display the results for Cases A and B (Case A/Case B in Table).
@@
In Table 3, we can observe the classification accuracy when we vary the data available to train the Vanilla MINT model for the best obtained Auditable Data (combination of all convolutional layers). Naturally, accuracy decreases when we reduce the number of training samples, demonstrating that the patterns to be learned are not trivial. A noteworthy result is that in the $50K$ scenario, accuracy decreases very slightly with respect to $100K$. This result suggests that an intermediate number of training samples may be enough to achieve acceptable results. Another interesting result is what occurs in the $1K$ scenario. For FR Model 1, performance decreases by $18\%$. However, for FR Model 2 and FR Model 3, it decreases by only $8\%$ and $6\%$, respectively. From these results, we can infer that depending on the model $M$ (its training, architecture, loss function, etc.), it may be possible to develop our MINT Model $T$ with more or less training data.
Table 3: Classification accuracy for the Vanilla MINT Model (Combination as Auditable Data) in the various data provided scenarios discussed for the three Face Recognition (FR) models.
@@
$100$K 0.84 0.84 0.81
$50$K 0.84 0.82 0.79
$1$K 0.66 0.76 0.75
In Table 4, we can observe the different classification accuracy results for the $3$ Face Recognition Models using the different information available to train the CNN MINT Model $T$ under the high data scenario. Here, the CNN is trained using the full activation blocks of each convolutional layer. Once again, the best performance is achieved with the Combination of all Conv Layers. Similarly, as observed previously, the next best results are achieved with Conv Layer #1 (i.e., the layer closest to the image domain), progressively decreasing as the layer gets closer to the end of the network (i.e., the layer closest to the output domain). The Model Outcome is not used, as explained in Sec. 3.3, because, as we discussed there, it is vectorial information that doesn’t fit in a CNN architecture in a straightforward way. For these CNN MINT Models, training with the combination of all the AADs significantly increases the computational complexity (approximately fourfold), while yielding only a modest improvement of around 1% in accuracy. Consequently, for the remaining experiments, we proceed with the CNN MINT Model trained using Conv Layer 1 as the AAD.
Table 4: Classification accuracy for various AAD using the CNN MINT model. MINT model trained with $100$K samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3).
@@
In Table 5, we observe the impact of varying the number of available data used to train the CNN MINT Model $T$ for the considered Auditable Data (Conv Layer #1). We see that for FR Model 1, a scenario with limited data for training the MINT Model negatively affect our results. However, for the other two FR Models, it has almost no effect on accuracy, achieving consistently high accuracy despite decreasing significantly the amount of training data. This indicates that for this MINT architecture depending on the model being audited, few samples ($500$ from $\mathcal{D}$) are enough to achieve good results. This result is very interesting as it implies that the auditor would require little cooperation from the developer. A similar situation occurs in Table 3 with the Vanilla MINT Model. The reason behind this is that, in Model 1, the activation patterns indicating whether a sample was used for training or not are more complex than in Models 2 and 3. As a result, the MINT Model requires more training data compared to the other two models to perform effectively.
Table 5: Classification accuracy for the CNN MINT Model (Conv Layer #1 as Auditable Data) in the various data provided scenarios discussed for the three Face Recognition (FR) models.
@@
$100$K 0.86 0.89 0.90
$50$K 0.85 0.89 0.87
$1$K 0.73 0.87 0.87
To conclude we compare our proposed MINT model to the state of the art, in Table 6. As mentioned earlier, to the best of our knowledge, there are no studies addressing the problem of detecting whether specific data was used for training state-of-the-art Face Recognition models. Therefore, making direct comparisons with other studies is not possible. Nevertheless, most studies around MIAs often use the classification output vector of the network [15, 38, 39, 17], or the output from the layers just before this output [40, 19]. These approaches are somewhat analogous to our Model Outcome (the Model Outcome of a Face Recognition Model is similar to the layers just before the output vector in a classification model), and thus, we use them as a reference for interpreting the state of the art, despite the environmental differences between the works around MIAs and our proposed MINT.
Table 6: Classification accuracy comparison between MIA and the two MINT approaches proposed in this work. We have adapted the MIA approach used in [19] using the Face Recognition Model outcome. Note that we used the original model instead of shadow models. ∗∗We used the Vanilla MINT model based on the combination of all convolutional layers. ∗∗∗We used the CNN MINT Model trained with the Conv Layer #1.
@@
Subsequently, we compare in Table 6 the best published MIA results with the best results achieved in our present study. As observed, the methods proposed in the present study significantly outperform what could be considered the state of the art in Face Recognition MIAs, surpassing it for all three FR Models by more than $10\%$. Finally, Fig. 5 provides a comparison of the ROC curves for the three methods outlined in the table 6, along with the results for the Vanilla network with information from individual Conv Layers. An important leap is evident between MIA (purple curve) and our best result using the CNN MINT Model (black curve), which we have also observed in Table 6. One of the main criticisms discussed in [19] and [16] is that the evaluation protocol and metrics commonly used in MIA approaches are not adequate, and the results presented in previous studies suffer from a high False Positive Rate. In Fig. 5, we can see that this is not the case in our study.
As explained in Sec. 4.2.3, an in-depth heuristic analysis (not fully reported here) has been conducted varying the complexity of the presented MINT Model architectures. From that study we can conclude that increasing the complexity of the architecture does not easily improve the performance. In our experiments, more complex architectures led to overfitting and a loss of performance in evaluation. To demonstrate this, Table 7 presents results by varying the number of parameters of the MINT CNN architecture. We provide results with the original model as explained in Sec. 4.2.3, another architecture with approximately three times the number of parameters, and the last one with ten times the number of parameters. We also include an experiment with a model three times simpler to demonstrate that the CNN requires enough complexity. Finally, we also present results using two additional architectures widely employed in the state of the art in computer vision. Specifically, the results obtained using a ResNet50 [76] and a ViT [77] as MINT Models.
Table 7: Classification accuracy comparison while varying the complexity of the MINT CNN network used. CNN MINT architecture with ÷$\div 3$ parameters. ∗∗ Original MINT CNN architecture presented in 4.2.3. ∗∗∗ CNN MINT architecture with ×$\times 3$ parameters. ∗∗∗∗ CNN MINT architecture with ×$\times 10$ parameters. ∗∗∗∗∗ ResNet50 [76] as CNN MINT architecture. ∗∗∗∗∗∗ ViT [77] as CNN MINT architecture
Method # Param. FR Model 1 FR Model 2 FR Model 3
@@
MINT CNN∗∗ $1.7$K 0.86 0.89 0.90
MINT CNN∗∗∗ $5$K 0.86 0.88 0.90
MINT CNN∗∗∗∗ $18$K 0.85 0.88 0.88
MINT ResNet50 $25.6$M 0.86 0.89 0.89
MINT ViT $86$M 0.86 0.89 0.90
6 Why Face Recognition? Exploring its Relevance and Broader Applications

### v4

@@ Abstract
Index Terms:
Audit, Fairness, Face Recognition, Membership Inference, MIA, MINT, Reliability.
@@
=-21pt
I Introduction
@@
Nowadays, AI systems have become increasingly advanced and pervasive across different domains, collecting, analyzing, and processing vast amounts of data. These data often contain sensitive information [3] about individuals as well as copyright content [4], raising concerns about privacy [5, 6, 7] and unauthorized access [8, 9, 10]. Consequently, this necessitates the fast development of novel tools to clarify what data has been used to train AI models, preventing therefore the use of non-authorized data and increasing the transparency and explainability of the AI models [11, 12, 13]. These tools aim to prevent learning frameworks from hiding the training data within the model’s parameters, thereby unveiling the opacity that they have traditionally relied upon. These considerations lead us to the main objective of the present study: checking if given data was used to train AI models [14], as summarized in Fig. 1.
@@
The article is structured as follows. Sec. II provides a revision of the state of the art in the scope of the article. Sec. III describes the main concepts of the proposed MINT, and the specific details for the task considered, Face Recognition. The databases, Face Recognition systems, and experimental protocol are described in Sec. IV. Results are discussed in Sec. V, whereas the significance of Face Recognition in MINT and the challenges of developing this technology in real-world applications are discussed in Sec. VI and VII. Discussion and conclusions are finally drawn in Sec. VIII and IX.
II Related Works
@@
II-A Membership Inference Attacks
@@
II-B Reconstructing Training Data from AI Models
@@
II-C Data Tracing
Methods described in previous sections can be seen as “passive methods” that rely on an already trained model without interfering with it. In contrast, the Data Tracing research direction falls under the category of “active methods”, where the training data is directly altered to influence the model [51]. Data tracing involves changing the training data to identify if it was used during training. Several studies have tried to tackle this challenge, demonstrating that with smart changes in the training data, it is feasible to detect its presence during inference [52, 53]. For instance, Radioactive Data [54] achieved this by inserting “radioactive marks” into the training data to detect if a database was used to train a model.
@@ II-C Data Tracing
III Membership Inference Test (MINT): Proposed Approach
@@
III-A Problem Statement and Terminology
@@
We hypothesize that an authorized auditor possesses access to model $M$, enabling him the acquisition of information regarding how $M$ processes data $d$. Alternatively, the auditor may possess information detailing how $M$ has processed data $d$, even in the absence of direct model access. This information comprises the generated Model Outcome $y=M(d|\textbf{w})$ and if possible also some intermediate results (e.g., activation maps of specific layers in a neural network). These intermediate outcomes $N(d|\textbf{w}^{\prime})$ provide insights about a subset of parameters $\textbf{w}^{\prime}\subset\textbf{w}$. We define these intermediate outcomes as Auxiliary Auditable Data (AAD). The auditor does not need the full description of the model $M$ or the values of the complete set of parameters w to obtain the AAD.
@@
  • Auxiliary Auditable Data $\textrm{AAD}=N(d|\textbf{w}^{\prime})$: intermediate outcomes of $M$ that result from processing an input data $d$ using a subset $\textbf{w}^{\prime}$ of the parameters w. The model outcome can be seen as the case where $\textbf{w}^{\prime}=\textbf{w}$ in which case $N(d|\textbf{w}^{\prime})=M(d|\textbf{w})$.
@@
III-B MINT: Application to Face Recognition Models
In the previous Sec. III-A, we have introduced the general concepts of our MINT proposal, which can be applied to any type of Audited Model $M$. Next, we detail the various components of the proposal for the application of MINT to the challenging task of Face Recognition.
@@ III-B MINT: Application to Face Recognition Models
  • $\mathcal{E}$ consists of images from external Face Recognition databases. We could add images unrelated to faces in $\mathcal{E}$. However, we included images with similar characteristics to $\mathcal{D}$ in order to prevent the network from learning to detect database-dependent features of face images instead of focusing on the main task: database-independent detection of used or not for training. It is crucial that these images do not overlap with those in $\mathcal{D}$, i.e., $\mathcal{E}\notin\mathcal{D}$.
@@
III-C Proposed MINT Models
@@ IV Experimental Framework
IV-A Face Recognition Data and Models
@@
As external data ($\mathcal{E}$) to train and test the MINT model $T$ we use the following datasets: The IJB-C [69] (3.5K identities and 138K face images), FDDB [70] (5.2K face images), GANDiffFace [71, 72] (10K identities and 500K face images), and Adience [73] (2.2K identities 26.5K face images). As mentioned in Sec. III-B, it is crucial to avoid overlap between $\mathcal{E}$ and $\mathcal{D}$, as this would lead to incorrect labels, compromising the training of MINT Models $T$. As noted by Schlett et al. [74], there is a significant number of duplicates across commonly used datasets. Although the data in $\mathcal{E}$ is sourced from evaluation datasets that strive to prevent such overlap, the necessary tests have been conducted to ensure that this issue does not occur.
IV-B Experimental Protocol
@@
  • 44 different databases as External Data $\mathcal{E}$. Three of them (IJB-C, FDDB, GANDiffFace) are used to extract samples for training the MINT Model, while the fourth one (Adience) is used for evaluation. We will refer to this choice of databases as the “Baseline Case”. Having either IJB-C or GANDiffFace, or both, in the training set is necessary to achieve balanced training (enough $\mathcal{E}$ samples). If we include only GANDiffFace from these two databases in the training set, the network might learn to recognize specific patterns from synthetic images instead of focusing on the MINT task. Therefore, we always include IJB-C in the training set. However, the remaining three databases are interchangeable, allowing evaluation with GANDiffFace or FDDB and utilization of Adience in training. We refer to these two cases as “Case A” and “Case B”, respectively. We present some results interchanging this evaluation database in our experiments. The Baseline case provides the most robust evaluation, as Case A involves evaluation with synthetic images, which may raise doubts due to the nature of these images, and Case B involves evaluation with only 5.2K images. Therefore, the majority of results will be presented under the “Baseline Case”.
@@ IV-B1 Data Pre-processing
IV-B2 Auxiliary Auditable Data and Model Outcome
@@
IV-B3 MINT Models Training
With the previously obtained AAD and Model Outcome, we propose two MINT models. Depending on how we handle the acquired information, we propose two architectures (Sec. III-C): a Vanilla MLP model and a CNN model. The CNN utilizes the activations of the convolutional blocks of $M$ as-is, while the Vanilla model, as previously explained, applies max pooling per channel to convert the activations of the convolutional blocks into a vector that can be processed by the MLP.
@@ IV-B3 MINT Models Training
Both architectures are chosen after extensive experimentation. It was found that the architectures did not need to be overly complex, as performance did not improve significantly with more complex architectures. In the experimental section (Section V-A), an experiment has been included to demonstrate the model’s performance versus the network complexity.
V Experiments and Results
This section presents the results achieved by our proposed MINT approach. As it involves binary classification between $\mathcal{D}$ and $\mathcal{E}$, and considers a balanced evaluation, the baseline accuracy (i.e., random guess) stands at 50%. Additionally, we recall that all results are derived from samples in $\mathcal{E}$ that do not belong to the $\mathcal{E}$ databases used in training the MINT Model, as elaborated in Sec. IV-B.
@@ V Experiments and Results
  • FR Model ($M$ in Fig. 3): These are the Face Recognition models presented in Sec. IV-A. The number corresponds to the one specified in that section.
  • Model Outcome: Refers to the Output Embedding of the Face Recognition model in Sec. III-B.
  • Conv Layer X: Represents the AAD obtained from each of the four primary layers, as explained in Sec. IV-B2
@@
  • Training Data: Represents three possible scenarios based on the amount of data available for training our MINT Model $T$. The high scenario (100100K) assumes a significant amount of data, with 5050K data $\mathcal{D}$ and 5050K external data $\mathcal{E}$. In the medium scenario (5050K), there are 2525K data samples each for $\mathcal{D}$ and $\mathcal{E}$. The low scenario (11K) considers only $500$ data samples for both $\mathcal{D}$ and $\mathcal{E}$. To put these numbers in context, 5050K data samples from $\mathcal{D}$ represent only $1\%$ of the MS1Mv3 database used for training and $0.3\%$ of Glint360k.
V-A Results
Table I shows the classification accuracy results for the $3$ Face Recognition models using different information to train the Vanilla MINT Model $T$ (AAD and Model Outcomes). All of this is done under a scenario of high data available for training. In these results, we can observe a superiority in accuracy when combining all the AAD information (Combination). In terms of the AAD it is important to highlight that in FR Model 1, Conv Layer #2 yields the best results after Conv Layer #1, while in FR Model 2 and 3, Conv Layer #4 follows with better accuracy. This might be attributed to the fact that ResNet architectures are based on “Skip Connections”, allowing information to traverse through the network from first layers to the last ones [76].
TABLE I: Classification accuracy for various AAD and Model Outcomes using the Vanilla MINT Model. MINT model trained with $100K$ samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3).
@@ V-A Results
The results in this section pertain to the Baseline Case presented in Sec. IV-B, where the databases IJB-C, FDDB, and GANDiffFace are used in training the MINT Models, while Adience is used for evaluation. This is the case that provides a more robust evaluation (Sec. IV-B), and therefore the one mostly used throughout the experimental section. However, to demonstrate the robustness of the results, Table II showcases the same results as Table I with exchanged training and evaluation databases for the other two scenarios (see Sec. IV-B):
@@
As shown in Table II, the outcomes are slightly superior for Case A and slightly inferior for Case B compared to the Baseline Case, yet the the conclusions remain unchanged (results are presented as Case A / Case B).
TABLE II: Classification accuracy for various AAD and Model Outcomes using the Vanilla MINT Model. MINT model trained with $100K$ samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3). We display the results for Cases A and B (Case A/Case B in Table).
@@
In Table III, we can observe the classification accuracy when we vary the data available to train the Vanilla MINT model for the best obtained Auditable Data (combination of all convolutional layers). Naturally, accuracy decreases when we reduce the number of training samples, demonstrating that the patterns to be learned are not trivial. A noteworthy result is that in the $50K$ scenario, accuracy decreases very slightly with respect to $100K$. This result suggests that an intermediate number of training samples may be enough to achieve acceptable results. Another interesting result is what occurs in the $1K$ scenario. For FR Model 1, performance decreases by $18\%$. However, for FR Model 2 and FR Model 3, it decreases by only $8\%$ and $6\%$, respectively. From these results, we can infer that depending on the model $M$ (its training, architecture, loss function, etc.), it may be possible to develop our MINT Model $T$ with more or less training data.
TABLE III: Classification accuracy for the Vanilla MINT Model (Combination as Auditable Data) in the various data provided scenarios discussed for the three Face Recognition (FR) models.
@@
100100K 0.84 0.84 0.81
5050K 0.84 0.82 0.79
11K 0.66 0.76 0.75
In Table IV, we can observe the different classification accuracy results for the $3$ Face Recognition Models using the different information available to train the CNN MINT Model $T$ under the high data scenario. Here, the CNN is trained using the full activation blocks of each convolutional layer. Once again, the best performance is achieved with the Combination of all Conv Layers. Similarly, as observed previously, the next best results are achieved with Conv Layer #1 (i.e., the layer closest to the image domain), progressively decreasing as the layer gets closer to the end of the network (i.e., the layer closest to the output domain). The Model Outcome is not used, as explained in Sec. III-C, because, as we discussed there, it is vectorial information that doesn’t fit in a CNN architecture in a straightforward way. For these CNN MINT Models, training with the combination of all the AADs significantly increases the computational complexity (approximately fourfold), while yielding only a modest improvement of around 1% in accuracy. Consequently, for the remaining experiments, we proceed with the CNN MINT Model trained using Conv Layer 1 as the AAD.
TABLE IV: Classification accuracy for various AAD using the CNN MINT model. MINT model trained with 100100K samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3).
@@
In Table V, we observe the impact of varying the number of available data used to train the CNN MINT Model $T$ for the considered Auditable Data (Conv Layer #1). We see that for FR Model 1, a scenario with limited data for training the MINT Model negatively affect our results. However, for the other two FR Models, it has almost no effect on accuracy, achieving consistently high accuracy despite decreasing significantly the amount of training data. This indicates that for this MINT architecture depending on the model being audited, few samples ($500$ from $\mathcal{D}$) are enough to achieve good results. This result is very interesting as it implies that the auditor would require little cooperation from the developer. A similar situation occurs in Table III with the Vanilla MINT Model. The reason behind this is that, in Model 1, the activation patterns indicating whether a sample was used for training or not are more complex than in Models 2 and 3. As a result, the MINT Model requires more training data compared to the other two models to perform effectively.
TABLE V: Classification accuracy for the CNN MINT Model (Conv Layer #1 as Auditable Data) in the various data provided scenarios discussed for the three Face Recognition (FR) models.
@@
100100K 0.86 0.89 0.90
5050K 0.85 0.89 0.87
11K 0.73 0.87 0.87
To conclude we compare our proposed MINT model to the state of the art, in Table VI. As mentioned earlier, to the best of our knowledge, there are no studies addressing the problem of detecting whether specific data was used for training state-of-the-art Face Recognition models. Therefore, making direct comparisons with other studies is not possible. Nevertheless, most studies around MIAs often use the classification output vector of the network [15, 38, 39, 17], or the output from the layers just before this output [40, 19]. These approaches are somewhat analogous to our Model Outcome (the Model Outcome of a Face Recognition Model is similar to the layers just before the output vector in a classification model), and thus, we use them as a reference for interpreting the state of the art, despite the environmental differences between the works around MIAs and our proposed MINT.
TABLE VI: Classification accuracy comparison between MIA and the two MINT approaches proposed in this work. We have adapted the MIA approach used in [19] using the Face Recognition Model outcome. Note that we used the original model instead of shadow models. ∗∗We used the Vanilla MINT model based on the combination of all convolutional layers. ∗∗∗We used the CNN MINT Model trained with the Conv Layer #1.
@@
Subsequently, we compare in Table VI the best published MIA results with the best results achieved in our present study. As observed, the methods proposed in the present study significantly outperform what could be considered the state of the art in Face Recognition MIAs, surpassing it for all three FR Models by more than $10\%$. Finally, Fig. 5 provides a comparison of the ROC curves for the three methods outlined in Table VI, along with the results for the Vanilla network with information from individual Conv Layers. An important leap is evident between MIA (purple curve) and our best result using the CNN MINT Model (black curve), which we have also observed in Table VI. One of the main criticisms discussed in [19] and [16] is that the evaluation protocol and metrics commonly used in MIA approaches are not adequate, and the results presented in previous studies suffer from a high False Positive Rate. In Fig. 5, we can see that this is not the case in our study.
As explained in Sec. IV-B3, an in-depth heuristic analysis (not fully reported here) has been conducted varying the complexity of the presented MINT Model architectures. From that study we can conclude that increasing the complexity of the architecture does not easily improve the performance. In our experiments, more complex architectures led to overfitting and a loss of performance in evaluation. To demonstrate this, Table VII presents results by varying the number of parameters of the MINT CNN architecture. We provide results with the original model as explained in Sec. IV-B3, another architecture with approximately three times the number of parameters, and the last one with ten times the number of parameters. We also include an experiment with a model three times simpler to demonstrate that the CNN requires enough complexity. Finally, we also present results using two additional architectures widely employed in the state of the art in computer vision. Specifically, the results obtained using a ResNet50 [76] and a ViT [77] as MINT Models.
TABLE VII: Classification accuracy comparison while varying the complexity of the MINT CNN network used. CNN MINT architecture with $\div 3$ parameters. ∗∗ Original MINT CNN architecture presented in IV-B3. ∗∗∗ CNN MINT architecture with $\times 3$ parameters. ∗∗∗∗ CNN MINT architecture with $\times 10$ parameters. ∗∗∗∗∗ ResNet50 [76] as CNN MINT architecture. ∗∗∗∗∗∗ ViT [77] as CNN MINT architecture
Method #Param FR 1 FR 2 FR 3
@@
MINT CNN∗∗ 1.71.7K 0.86 0.89 0.90
MINT CNN∗∗∗ 55K 0.86 0.88 0.90
MINT CNN∗∗∗∗ 1818K 0.85 0.88 0.88
MINT ResNet50 25.625.6M 0.86 0.89 0.89
MINT ViT 8686M 0.86 0.89 0.90
VI Why Face Recognition? Exploring its Relevance and Broader Applications